that a firms recent earnings per share and dividend per share are $3.30 and $2.90, respectively Both are expected to grow at 6 percent. However, the firms current P/E ratio of 30 seems high for this growth rate. The P/E ratio is expected to fall to 26 within five years. Compute the dividends over the next five years. (Do not round intermediate calculations. Round your final answer to 3 decimal places.) Years First year Second year Third year Fourth year Fifth year Compute the value of this stock price in five years. (Do not round intermediate calculations. Round your final answer to 2 decimal places.) 38.86 Stock price Calculate the present value of these cash flows using an 8 percent discount rate. (Do not round intermediate calculations. Round your final answer to 2 decimal places.) Present value
Answer #1

The dividend should be as follows:
Year 1 = 2.90*1.06=$3.074

The complete table is as below:

Dividend EPS
First year $    3.074 $   3.498
Second $    3.258 $   3.708
Third $    3.454 $   3.930
Fourth $    3.661 $   4.166
Fifth $    3.881 $   4.416

Now the PE ratio drops to 26, Hence year end price = 4.416*26=$114.82

The present value of all cash flows =

Year Dividend Terminal value Total DF PV
1 $    3.074 $       3.07 0.925926 $       2.85
2 $    3.258 $       3.26 0.857339 $       2.79
3 $    3.454 $       3.45 0.793832 $       2.74
4 $    3.661 $       3.66 0.73503 $       2.69
5 $    3.881 $ 114.82 $ 118.70 0.680583 $    80.79
Total $    91.86

DF is calculated as 1/(1.08^n) where n is the number of year.

The present value of cash flow is therefore $91.86